OpenStudy (anonymous):

how do i find the six trig functions with one point being (-4,0) and its in standard position?

OpenStudy (mathstudent55):

You need to know the definitions of all the 6 functions. \(\sin \alpha = \dfrac{x}{r} \) \(\cos \alpha = \dfrac{y}{r} \) \(\tan \alpha = \dfrac{y}{x} \) \(\sec \alpha = \dfrac{r}{x} \) \(\csc \alpha = \dfrac{r}{y} \) \(\tan \alpha = \dfrac{y}{x} \)